Caddie Manufacturing has a target debt-equity ratio of .60. Its cost of equity is 11 percent, and its pretax cost of debt is 6 percent. If the tax rate is 22 percent, what is the company’s WACC?

WACC %
Answer #1

After-tax cost of debt=6*(1-tax rate)

=6*(1-0.22)=4.68%

debt-equity ratio=debt/equity

Hence debt=0.6*equity

Let equity be $x

Debt=$0.6x

Total=$1.6x

WACC=Respective costs*Respective weight

=(x/1.6x*11)+(0.6x/1.6x*4.68)

=8.63%